OpenStudy (anonymous):

Rewrite the following equation from point-slope form to slope-intercept form.

OpenStudy (anonymous):

y-1=4(x+2)

OpenStudy (pfenn1):

rearrange to solve for y in the form of y=mx+b

OpenStudy (anonymous):

how would i rearange it

OpenStudy (lgbasallote):

first, distribute 4 into your quantity..

OpenStudy (anonymous):

what does that mean

OpenStudy (pfenn1):

y-1=4x+8 y-1+1=4x+8+1 y=4x+9